# Near-linear Time Dispersion of Mobile Agents

### v1

@@ Abstract
Consider that there are $k\leq n$ agents in a simple, connected, and undirected graph $G=(V,E)$ with $n$ nodes and $m$ edges. The goal of the dispersion problem is to move these $k$ agents to distinct nodes. Agents can communicate only when they are at the same node, and no other means of communication such as whiteboards are available. We assume that the agents operate synchronously. We consider two scenarios: when all agents are initially located at any single node (rooted setting) and when they are initially distributed over any one or more nodes (general setting). Kshemkalyani and Sharma presented a dispersion algorithm for the general setting, which uses $O(m_{k})$ time and $\log(k+\delta)$ bits of memory per agent [OPODIS 2021]. Here, $m_{k}$ is the maximum number of edges in any induced subgraph of $G$ with $k$ nodes, and $\delta$ is the maximum degree of $G$. This algorithm is the fastest in the literature, as no algorithm with $o(m_{k})$ time has been discovered even for the rooted setting. In this paper, we present faster algorithms for both the rooted and general settings. First, we present an algorithm for the rooted setting that solves the dispersion problem in $O(k\log\min(k,\delta))=O(k\log k)$ time using $O(\log\delta)$ bits of memory per agent. Next, we propose an algorithm for the general setting that achieves dispersion in $O(k(\log k)\cdot(\log\min(k,\delta))=O(k\log^{2}k)$ time using $O(\log(k+\delta))$ bits.
@@ 2 Preliminaries
We evaluate the time complexity or running time of algorithm $\mathcal{A}$ as the maximum number of steps until $\Xi_{\mathcal{A}}(C_{0})$ reaches a legitimate configuration, where the maximum is taken over all initial configurations $C_{0}$. Let $\mathcal{M}^{\prime}_{\mathcal{A}}\subseteq\mathcal{M}_{\mathcal{A}}$ be the set of all agent-states that can appear in any possible execution of $\mathcal{A}$ starting from any initial configuration. We evaluate the space complexity or memory space of algorithm $\mathcal{A}$ as $\log_{2}|\mathcal{M}^{\prime}_{\mathcal{A}}|$, i.e., the maximum number of bits required to represent an agent-state that may appear in those exeuctions. This implies that we exclude the size of the working memory used for deciding the destination and updating states, and the size of the space for storing input information. In particular, the space complexity can be $o(\log k)$ or $o(\log\delta)$, while each agent reads its identifier and the incoming port at each time step.222 We adopt this definition following the convention in the study of mobile agents. Some papers dealing with the dispersion problem count the number of bits required for storing the agent identifier and the incoming port.
Throughout this paper, we denote by $[i,j]$ the set of integers $\{i,i+1,...,j\}$. We have $[i,j]=\emptyset$ when $j<i$. We frequently use $\tau=\min(k,\delta)$. We define $\nu(a,t)$ as the node where agent $a$ resides at time step $t$. We also omit time step $t$ from any function in the form $f(*,t)$ and just write $f(*)$ if $t$ is clear from the context. For example, we just write $A(v)$ and $\nu(a)$ instead of $A(v,t)$ and $\nu(a,t)$.
@@ 3 Rooted Dispersion
1 $b.\mathtt{settled}\leftarrow\top$, where $b$ is the agent with the smallest ID in $A(a_{\mathrm{max}})$
@@
4             All explorers in $A(w)$ goes to $u$ $b^{\prime}.\mathtt{settled}\leftarrow\top$, where $b^{\prime}$ is the agent with the smallest ID in $A(u)$ $b^{\prime}.\mathtt{parent}\leftarrow a_{\mathrm{max}}.\mathtt{pin}$
@@
6             All explorers in $A(w)$ goes back to node $N(w,\psi(w).\mathtt{parent})$.
@@ 4.1 Overview
Unlike $\mathbf{RootedDispersion}$, a leader does not settle itself at a node in the final stage of HEO-DFS. The leader $a$ suspends the HEO-DFS if it visits an unsettled node but it has no accompanying zombies to settle at that time. A leader who has suspended the HEO-DFS due to the absence of accompanying zombies is called a waiting leader. Conversely, a leader with accompanying zombies is called an active leader. A waiting leader $a$ resumes the HEO-DFS when a zombie cathces up to $a$ at $\nu(a)$. As we will see later, the execution of $\mathbf{GeneralDispersion}$ ensures that all agents eventually become either waiting leaders or settlers, each residing at a distinct node. Since the dispersion problem does not require termination detection, the agents have solved the dispersion once such a configuration is reached.
@@ 7                         Settle one zombie in $A_{Z}(w)$ at $w$
24             $a_{i}$ moves to $u_{i}$. /*************** Slot 7 begins ***************/ if $(a.\mathtt{leader},a.\mathtt{level})=(\psi(u_{i}).\mathtt{leader},\psi(u_{i}).\mathtt{level})$ then
@@ 4.2 Details
Note that, even during the probing process at node $w$, leader $a$ might become a zombie if it meets a stronger leader $b$ in slot 1. Some settlers might then move to $w$ in the next slot 3 to help $a$, not knowing $a$ is now a zombie. In these situations, $b$ changes the $\mathtt{help}$ of these settlers to $\bot$ and sends them back to their homes in slot 4. Thereafter, those setllers remain at their home at least until they are incoporated into another group.

### v2

@@ Abstract
Consider that there are $k\leq n$ agents in a simple, connected, and undirected graph $G=(V,E)$ with $n$ nodes and $m$ edges. The goal of the dispersion problem is to move these $k$ agents to distinct nodes. Agents can communicate only when they are at the same node, and no other means of communication such as whiteboards are available. We assume that the agents operate synchronously. We consider two scenarios: when all agents are initially located at any single node (rooted setting) and when they are initially distributed over any one or more nodes (general setting). Kshemkalyani and Sharma presented a dispersion algorithm for the general setting, which uses $O(m_{k})$ time and $\log(k+\delta)$ bits of memory per agent [OPODIS 2021]. Here, $m_{k}$ is the maximum number of edges in any induced subgraph of $G$ with $k$ nodes, and $\delta$ is the maximum degree of $G$. This algorithm is the fastest in the literature, as no algorithm with $o(m_{k})$ time has been discovered even for the rooted setting. In this paper, we present faster algorithms for both the rooted and general settings. First, we present an algorithm for the rooted setting that solves the dispersion problem in $O(k\log\min(k,\delta))=O(k\log k)$ time using $O(\log\delta)$ bits of memory per agent. Next, we propose an algorithm for the general setting that achieves dispersion in $O(k(\log k)\cdot(\log\min(k,\delta))=O(k\log^{2}k)$ time using $O(\log(k+\delta))$ bits. Finally, for the rooted setting, we give a time-optimal, i.e., $O(k)$-time, algorithm with $O(\delta)$ bits of space per agent.
@@ 1.1 Our Contribution
While the two algorithms mentioned above are nearly time-optimal, i.e., requiring $O(k\cdot\log^{c}k)$ time for some constant $c$, they are not entirely optimal. We also demonstrate that in the rooted setting, a time-optimal algorithm based on the HEO-DFS can be achieved if significantly more space is available, specifically $O(\delta)$ bits per agent.
@@ 2 Preliminaries
We evaluate the time complexity or running time of algorithm $\mathcal{A}$ as the maximum number of steps until $\Xi_{\mathcal{A}}(C_{0})$ reaches a legitimate configuration, where the maximum is taken over all initial configurations $C_{0}$. Let $\mathcal{M}^{\prime}_{\mathcal{A}}\subseteq\mathcal{M}_{\mathcal{A}}$ be the set of all agent-states that can appear in any possible execution of $\mathcal{A}$ starting from any initial configuration. We evaluate the space complexity or memory space of algorithm $\mathcal{A}$ as $\log_{2}|\mathcal{M}^{\prime}_{\mathcal{A}}|$, i.e., the maximum number of bits required to represent an agent-state that may appear in those executions. This implies that we exclude the size of the working memory used for deciding the destination and updating states, and the size of the space for storing input information. In particular, the space complexity can be $o(\log k)$ or $o(\log\delta)$, while each agent reads its identifier and the incoming port at each time step.222 We adopt this definition following the convention in the study of mobile agents. Some papers dealing with the dispersion problem count the number of bits required for storing the agent identifier and the incoming port.
Throughout this paper, we denote by $[i,j]$ the set of integers $\{i,i+1,...,j\}$. We have $[i,j]=\emptyset$ when $j<i$. When the base of a logarithm is not specified, it is assumed to be 2. We frequently use $\tau=\min(k,\delta)$. We define $\nu(a,t)$ as the node where agent $a$ resides at time step $t$. We also omit time step $t$ from any function in the form $f(*,t)$ and just write $f(*)$ if $t$ is clear from the context. For example, we just write $A(v)$ and $\nu(a)$ instead of $A(v,t)$ and $\nu(a,t)$.
@@ 3 Rooted Dispersion
1 $b.\mathtt{settled}\leftarrow\top$, where $b$ is the agent with the smallest ID in $A(s)$
@@
4             All explorers in $A(w)$ go to $u$ $b^{\prime}.\mathtt{settled}\leftarrow\top$, where $b^{\prime}$ is the agent with the smallest ID in $A(u)$ $b^{\prime}.\mathtt{parent}\leftarrow a_{\mathrm{max}}.\mathtt{pin}$
@@
6             All explorers in $A(w)$ go back to node $N(w,\psi(w).\mathtt{parent})$.
@@ 4.1 Overview
Unlike $\mathbf{RootedDispersion}$, a leader does not settle itself at a node in the final stage of HEO-DFS. The leader $a$ suspends the HEO-DFS if it visits an unsettled node but it has no accompanying zombies to settle at that time. A leader who has suspended the HEO-DFS due to the absence of accompanying zombies is called a waiting leader. Conversely, a leader with accompanying zombies is called an active leader. A waiting leader $a$ resumes the HEO-DFS when a zombie catches up to $a$ at $\nu(a)$. As we will see later, the execution of $\mathbf{GeneralDispersion}$ ensures that all agents eventually become either waiting leaders or settlers, each residing at a distinct node. Since the dispersion problem does not require termination detection, the agents have solved the dispersion once such a configuration is reached.
@@ 7                         Settle one zombie in $A_{Z}(w)$ at $w$
24             $a_{i}$ moves to $u_{i}$. /*************** Slot 7 begins ***************/ if $(a_{i}.\mathtt{leader},a_{i}.\mathtt{level})=(\psi(u_{i}).\mathtt{leader},\psi(u_{i}).\mathtt{level})$ then
@@ 4.2 Details
Note that, even during the probing process at node $w$, leader $a$ might become a zombie if it meets a stronger leader $b$ in slot 1. Some settlers might then move to $w$ in the next slot 3 to help $a$, not knowing $a$ is now a zombie. In these situations, $b$ changes the $\mathtt{help}$ of these settlers to $\bot$ and sends them back to their homes in slot 4. Thereafter, those settlers remain at their home at least until they are incorporated into another group.
@@
5 For Further Improvement in Time Complexities
1 $\mathtt{Settle}(b,s,\bot)$, where $s$ is the starting node and $b$ is the agent with the smallest ID in $A(s)$ $\mathtt{Probe}()$ while $a_{\mathrm{max}}.\mathtt{settled}=\bot$ do
2       Let $w=\nu(a_{\mathrm{max}})$ if $U_{0}(w)\neq\bot$ then
             Chose any neighbor $u\in U_{0}(w)$   // $u$ is an unsetteled node here
3             All explorers in $A(w)$ go to $u$ $\mathtt{Settle}(b,u,a_{\mathrm{max}}.\mathtt{pin})$, where $b^{\prime}$ is the agent with the smallest ID in $A(u)$ $\mathtt{Probe}()$
4       else
5             All explorers in $A(w)$ go back to node $N(w,\psi(w).\mathtt{parent})$.
6
7function $\mathtt{Settle}(b,v,p)$:
8       $b.\mathtt{settled}\leftarrow\top$ $b.\mathtt{parent}\leftarrow p$ for $i\in[0,\delta_{v}]$ do
9            $b.\mathtt{checked}[i]\leftarrow\bot$
10
11function $\mathtt{Probe}()$:
       Let $w=\nu(a_{\mathrm{max}})$ Let $\ell=|A(w)\setminus\{\psi(w)\}|$   // $\ell$ is the number of explorers
12       while $U_{\bot}(w)\neq\emptyset\land|U_{0}(w)|<\ell$ do
             Let $a_{1},a_{2},\dots,a_{x}$ be the agents in $A(w)\setminus\{\psi(w)\}$  // $x\leq\ell$
13             Let $\delta^{\prime}=\min(x,|U_{\bot}(w)|)$ Choose $\delta^{\prime}$ neighbors $u_{1},u_{2},\dots,u_{\delta^{\prime}}$ from $U_{\bot}$ arbitrarily. for each $a_{i}\in\{a_{1},a_{2},\dots,a_{\delta^{\prime}}\}$ in parallel do
14                   $a_{i}$ moves to $u_{i}$ if $a_{i}$ finds $\psi(u_{i})$ at $u_{i}$ then
15                         $\psi(u_{i}).\mathtt{checked}[a_{i}.\mathtt{pin}]\leftarrow 1$
16                  $a_{i}$ goes back to $w$ along with $\psi(u_{i})$ (if it exists) Let $p_{i}$ be a port number such that $N(w,p_{i})=u_{i}$ $\psi(w).\mathtt{checked}[p_{i}]\leftarrow\begin{cases}1&\text{if }a_{i}\text{ found }\psi(u_{i})\text{ at }u_{i}\\ 0&\text{otherwise},\end{cases}$
17
18      Let all settlers except for $\psi(w)$ go back to their homes if $|U_{0}(w)|\geq\ell$ then
19             All explorers in $A(w)$ go to distinct neighbors in $U_{0}(w)$ and settle at those nodes.
20
Algorithm 6 $\mathbf{RootedOptimal}$
In the previous sections, we introduced nearly time-optimal dispersion algorithms: an $O(k\log\tau)$-time algorithm for the rooted setting and an $O(k\log^{2}k)$-time algorithm for the general setting. This raises a crucial question: is it possible to develop a truly time-optimal algorithm, specifically an $O(k)$-time algorithm, even if it requires much more space? In this section, we affirmatively answer this question for the rooted setting. We present an $O(k)$-time algorithm that utilizes $O(\delta)$ bits of space per agent. However, the feasibility of an $O(k)$-time algorithm in the general setting remains open.
We refer to the new algorithm as $\mathbf{RootedOptimal}$ in this section. In $\mathbf{RootedDispersion}$, with $O(\log\delta)$ bits of space, each settler $\psi(w)$ cannot memorize the exact set of settled neighbors of $w$. Instead, it only remembers that the first $i$ neighbors of $w$ are settled for some $i$. In contrast, $\mathbf{RootedOptimal}$ allows each settler $\psi(w)$ to remember all settled neighbors of $w$ using $O(\delta)$ bits, which significantly helps to eliminate an $O(\log\tau)$ factor from the time complexity. However, somewhat surprisingly, both the design of the new algorithm and the analysis of its execution time are non-trivial.
Below, we outline the modifications made to $\mathbf{RootedDispersion}$ to obtain $\mathbf{RootedOptimal}$. The pseudocode for $\mathbf{RootedOptimal}$ is provided in Algorithm 6. However, we anticipate that most readers will grasp the behavior of $\mathbf{RootedOptimal}$ simply by reviewing the following key differences.
  • In $\mathbf{RootedOptimal}$, each settler $\psi(w)$ maintains an array variable $\psi(w).\mathtt{checked}$ of size $\delta_{w}$. Each element of $\psi(w).\mathtt{checked}[i]$ takes a value from the set $\{0,1,\bot\}$. The assignment $\psi(w).\mathtt{checked}[i]=0$ (respectively, $1$) indicates that the neighbor $N(w,i)$ is unsettled (respectively, settled). The value $\bot$ is utilized exclusively during the probing process, meaning that the neighbor $N(w,i)$ has yet to be checked for its settled status. For any $c\in\{0,1,\bot\}$, we define $U_{c}=\{N(u,i)\mid i\in[0,\delta_{w}-1],\psi(w).\mathtt{checked}[i]=c\}$.
  • Consider that the unique leader $a_{\mathrm{max}}$ invokes the probing process $\mathtt{Probe}()$ at a node $w$. (Remember that $a_{\mathrm{max}}$ is the unique leader that has the maximum identifier at the beginning of the execution.) In $\mathbf{RootedDispersion}$, the probing process terminates as soon as any agent finds an unsettled neighbor. However, in $\mathbf{RootedOptimal}$, the process only ends when all of $w$’s neighbors are probed or when at least $\ell$ unsettled neighbors are found, where $\ell$ is the number of explorers. In the former case, the probing process is now complete: $U_{\bot}(w)$ is empty, and $U_{0}(w)$ equals the set of unsettled neighbors of $w$. In the latter case, the explorers go to distinct $\ell$ unsettled nodes and settle there, thereby achieving dispersion.
  • Consider an agent $a$ making a round trip $w\to u\to w$ during $\mathtt{Probe}()$, where $u=N(w,p)$ for some $p\in[0,\delta_{w}-1]$. If $a$ does not encounter a settler at $u$, it simply sets $\psi(w).\mathtt{checked}[i]$ to 00. On the other hand, if a settler is found at $u$, $a$ sets $\psi(w).\mathtt{checked}[i]$ to $1$ and additionally sets $\psi(u).\mathtt{checked}[q]$ to $1$, where $q\in[0,\delta_{u}-1]$ is the port number such that $w=N(u,q)$. This modification ensures that $U_{0}(u)$ remains equal to the set of unsettled neighbors of $u$ when $\mathtt{Probe}$ finishes.
  • In $\mathbf{RootedDispersion}$, the leader $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ after each forward or backward move. However, in $\mathbf{RootedOptimal}$, $a_{\mathrm{max}}$ only invokes $\mathtt{Probe}()$ after making a forward move. This change does not compromise the correctness of $\mathbf{RootedOptimal}$ because when $a_{\mathrm{max}}$ makes a backward move to a node $w$, $\psi(w)$ accurately remembers its unsettled neighbors due to the modification mentioned earlier.
One might think that the probing process $\mathtt{Probe}()$ in $\mathbf{RootedOptimal}$ could take longer time than in $\mathbf{RootedDispersion}$, as it only finishes after all neighbors of the current location have been probed or after finding $\ell$ unsettled neighbors, where $\ell$ is the number of explorers. Particularly, there seems to be a concern that during the probing at a node $w$, the number of agents, excluding $\psi(w)$, may not always double: this event occurs when some agents discover an unsettled neighbor. Despite that, we deny this conjecture at least asymptotically, that is, we have the following lemma.
Lemma 7.
Assume that $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at node $w$ during the execution of $\mathbf{RootedOptimal}$, and exactly $\ell$ explorers including $a_{\mathrm{max}}$ exists at the time. Then, $\mathtt{Probe}()$ finishes within $O(1)+\max(0,2\lceil\log\tau-\log\ell\rceil)$ time.
Proof.
If $\ell\geq\tau$, we have $\ell\geq\min(\delta,k)=\delta$ because $\ell<k$. Then, the lemma trivially holds: $\mathtt{Probe}()$ finishes in a constant time. Thus, we consider the case $\ell<\tau$. Let $t$ be the time step at which $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at a node $w$, and let $z=\lceil\log\tau-\log\ell\rceil+1$. It suffices to show that $U_{\bot}(w,t^{\prime})=\emptyset$ or $U_{0}(w,t^{\prime})\geq\ell$ holds for some $t^{\prime}\in[t,t+2z]$. Assume for contradiction that this does not hold. For any $r\in[0,z]$, we define
$f(r)=X_{r}\cdot 2^{r-1}+Y_{r},$
where $X_{r}=|U_{0}(w,t+2r)|$ and $Y_{r}=|A(w,t+2r)\setminus\{\psi(w)\}|$. By definition, $f(0)=0\cdot 2^{0-1}+\ell=\ell$. Under the above assumption, for any $r=0,1,\dots,z-1$, the agents in $A(w,t+2r)$ move to distinct neighbors in $U_{\bot}(w,t+2r)$ in time step $t+2r$, and bring back all settlers they find, at most one for each neighbor, in time step $t+2r+1$. Let $\alpha$ be the number of those settlers i.e.$\alpha=Y_{r+1}-Y_{r}$. Note that $X_{r+1}-X_{r}=Y_{r}-\alpha$ hods here. Then, irrespective of $\alpha$, we obtain
$f(r+1)$ $=X_{r+1}\cdot 2^{r}+Y_{r+1}$
$=(X_{r}+Y_{r}-\alpha)2^{r}+Y_{r}+\alpha$
$=X_{r}\cdot 2^{r}+(2^{r}-1)(Y_{r}-\alpha)+2\cdot Y_{r}$
$\geq 2(x_{r}\cdot 2^{r-1}+Y_{r})$
$=2f(r),$
where we use $2^{r}\geq 1$ and $Y_{r}\geq\alpha$ in the above inequality. Therefore, we have $f(z)\geq\ell\cdot 2^{z}$, while $f(z)=X_{z}\cdot 2^{z-1}+Y_{z}\leq(\ell-1)2^{z-1}+Y_{z}$ by the assumption that $X_{z}=|U_{0}(w,t+2z)|<\ell$. Thus, we have $|A(w,t+2z)\setminus\{\psi(w)\}|=Y_{z}\geq\ell\cdot 2^{z}-(\ell-1)2^{z-1}=(\ell+1)2^{z-1}\geq(\ell+1)\tau/\ell>\tau$. Since $\tau=\min(\delta,k)$, we have $\tau=\delta$ or $\tau=k$. In the former case, $Y_{k}>\delta$ agents at $w$ are enough to visit all neighbors in $U_{\bot}(w,t+2z)$ in time step $t+2z$, thus $U_{\bot}(w,t+2z+2)=\emptyset$ holds, a contradiction. In the latter case, there are $|A(w,t+2z)|\geq k+2$ agents in $w$ at time step $t+2z$, a contradiction. Therefore, $U_{\bot}(w)=\emptyset$ or $U_{0}(w)\geq\ell$ holds at time step $t^{\prime}\leq t+2z+2$. ∎
Theorem 3.
In the rooted setting, algorithm $\mathbf{RootedOptimal}$ solves the dispersion problem within $O(k)$ time using $O(\delta)$ bits of space per agent.
Proof.
The unique leader $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ only when it settles an agent, except for when $a_{\mathrm{max}}$ itself becomes settled. Therefore, $a_{\mathrm{max}}$ invokes $\mathtt{Probe}$ exactly $k-1$ times, with precisely $k-i$ explorers present at the $i$-th invocation. By Lemma 7, the total number of steps required for the $k-1$ executions of $\mathtt{Probe}()$ is at most
$\sum_{\ell=1}^{k-1}(\log k-\log\ell)+O(k)\leq k\log k-\log(k!)+\log k+O(k)=O(k),$
where we apply Stirling’s formula, i.e.$\log(k!)=k\log k-k+O(\log k)$. As demonstrated in Section 3, both forward and backward moves also require a total time of $O(k)$. Thus, $\mathbf{RootedOptimal}$ completes in $O(k)$ time. Regarding space complexity, the array variable $\mathtt{checked}$ is the primary factor, needing $O(\delta)$ bits per agent. Other variables require only $O(\log\delta)$ bits. ∎